OpenStudy (anonymous):

Derive: sin2x=2sinxcosx

OpenStudy (anonymous):

You can take the derivative of either side (though the lefthand side will be easier).

OpenStudy (anonymous):

sin(2x)=sin(x+x) =sinxcosx+cosxsinx FORMULA :(sin(A+B)=sinAcosB+cosAsinB) =2sinxcosx

OpenStudy (anonymous):

considering how the derivative of sinx is cosx i believe that with assumption that \[\lim_{x \rightarrow 2x}\] ( it is valid ) the derivative of sin2x will be cos2x which is equal to \[1-2 \sin ^{2} x\] This is however simply my assumption